# Measuring Data Science Automation: A Survey of Evaluation Tools for AI Assistants and Agents

### v1

@@ 4.3 End-to-end tasks scored by their final result
The tasks require touch upon multiple data science activities—from Goal Exploration to Data Understanding, Value Exploration and Preparation, and to Modelling and Narrative Exploration—but only the final insights are directly evaluated (Deployment, Narrative Exploration, and Result Exploration). Similarly, BLADE (Gu et al., 2024) and DiscoveryBench (Majumder et al., 2024) both challenge LLM agents to explore a complex dataset with a vaguely defined goal, such as “Are soccer players with a dark skin tone more likely than those with a light skin tone to receive red cards from referees?” (Gu et al., 2024). However, differently from InsightBench, they consider scientific datasets and focuses on agents’ ability to integrate statistical knowledge with understanding of data from a broad range of scientific domains. Both BLADE and DiscoveryBench include end-to-end scientific data-analysis tasks that begin with genuine research questions and necessitate multistep solution workflows. BLADE includes 12 carefully curated datasets collected from statistical textbooks, research papers, and crowd-sourced studies, while DiscoveryBench offers 264 real-world tasks drawn from published studies plus 903 synthetic tasks spanning 48 domains. Tasks cover Data Understanding, Preparation, statistical and Machine Learning Modelling, Narrative Exploration, as well as Deployment, Data Value Exploration and various degrees of domain understanding (Business Understanding and Data Understanding). Both BLADE and DiscoveryBench grade solutions automatically with LLMs so that multiple defensible workflows can receive credit. Their emphases, however, diverge: BLADE checks if each analytical step—conceptual variable selection, admissible transformations, model family, hyper-parameters—corresponds to one of multiple expert-produced solutions (to account for alternatives), which however still limits the amount by which the agent can transform the task; DiscoveryBench instead scores at the level of the final context, variables and relationship identified, using GPT-4 to judge the semantic match between the agent’s claim and without considering how the former was obtained. An overview of BLADE and DiscoveryBench can be found in Figs. 2 and 5.
@@
From Table 3, we can see that agents have been evaluated on more data science activities than assistants (Table 2), particularly considering goal-oriented activities (with the exception of Business Understanding); however, there is still a lack of evaluations for data management and, to a lesser extent, exploratory activities. Additionally, of all the surveyed works, only BiasBenchmark (Li et al., 2025b), Tapilot-Crossing (Li et al., 2024), and IDA-Bench (Li et al., 2025a) evaluate agents in a collaborative framework with (simulated) users.
@@ 5 Challenges and future directions
Our analysis shows that most evaluation works focus either on assistance, looking at isolated tasks that require LLMs to provide an answer on a single-turn basis (without access to tools and under human supervision) or on full automation, where LLMs are wrapped in agents that act autonomously. A few notable exceptions exist (Yu et al., 2019a; Li et al., 2024; 2025b; 2025a), which primarily rely on other LLMs to simulate human users. While this approach ensures cost-effective evaluation and reproducibility, exploratory tasks may lead the agent to attempt novel solutions, such that the simulated user may be unable to assist it as the most suitable answer may not be within its knowledge base. This can be partly addressed by employing humans to answer such unprecedented queries and progressively enriching the knowledge base. Furthermore, informative evaluations should also quantify the trade-off between autonomy and reliability/performance; among the surveyed works, only Li et al. (2025a) advances in this direction.
We found data management and exploratory activities remain mostly uncovered. This is due to 1) the inherent difficulty of scoring exploratory activities, which lack a fixed ground truth, and 2) the complex real-world interactions that certain exploratory activities (such as Business Understanding and Goal Exploration) and data management activities (Acquisition and Simulation) demand. To address this issue, simulated environments where data management and client interaction can occur should be developed, analogous to related developments in scientific research evaluation (Jansen et al., 2024; Cerrato et al., 2024). Such environments would enable the evaluation of agents or assistants that function holistically as data scientists by understanding business requirements, facilitating data collection, and adapting customer requests through data exploration. Simultaneously, realistic evaluation should progress toward end-to-end tasks that do not depend on strict ground truth comparisons or simple activity-specific metrics. Instead, these evaluations should reward insight generation (such as the works in Sec. 4.3) in potentially original ways, thereby properly incentivising systems that fundamentally redefine activities rather than focusing solely on human substitution.
Thus, we propose the following actions to improve data science evaluation of AI systems:
@@
Acknowledgments
IT and LP received funding from Open Philanthropy. JHO acknowledges CIPROM/2022/6 (FASSLOW) funded by Generalitat Valenciana, and Spanish grant PID2021-122830OB-C42 (SFERA) funded by MCIN/AEI/10.13039/501100011033 and “ERDF A way of making Europe" Cátedra ENIA-UPV in Sustainable AI Development, TSI-100930-2023-9, and INCIBE’s Chair funded by the EU-NextGenerationEU through the Spanish government’s Plan de Recuperación, Transformación y Resiliencia, and EUR2024-153548 (PREDAIT) "Towards Predictable AI" from Spanish "Europa Excelencia" 2024.
@@ Appendix B Appendix: examples of tasks
Figure 4: Li et al. (2024): The figure shows the “Normal” mode, with the agent being provided all the relevant information and tasked with writing code to address the task, and “Action” mode, where the agent has to take a specific action (in this case, asking for clarification). “Private” refers to tasks requiring the use of bespoke software libraries to which the agent has access to.

### v2

@@ 4.3 End-to-end tasks scored by their final result
The tasks require touch upon multiple data science activities—from Goal Exploration to Data Understanding, Value Exploration and Preparation, and to Modelling and Narrative Exploration—but only the final insights are directly evaluated (Deployment, Narrative Exploration, and Result Exploration). Similarly, BLADE (Gu et al., 2024) and DiscoveryBench (Majumder et al., 2024) both challenge LLM agents to explore a complex dataset with a vaguely defined goal, such as “Are soccer players with a dark skin tone more likely than those with a light skin tone to receive red cards from referees?” (Gu et al., 2024). However, differently from InsightBench, they consider scientific datasets and focus on agents’ ability to integrate statistical knowledge with understanding of data from a broad range of scientific domains. Both BLADE and DiscoveryBench include end-to-end scientific data‐analysis tasks that begin with genuine research questions and necessitate multistep solution workflows. BLADE includes 12 carefully curated datasets collected from statistical textbooks, research papers, and crowd-sourced studies, while DiscoveryBench offers 264 real-world tasks drawn from published studies plus 903 synthetic tasks spanning 48 domains. Tasks cover Data Understanding, Preparation, statistical and Machine Learning Modelling, Narrative Exploration, as well as Deployment, Data Value Exploration and various degrees of domain understanding (Business Understanding and Data Understanding). Both BLADE and DiscoveryBench grade solutions automatically with LLMs so that multiple defensible workflows can receive credit. Their emphases, however, diverge: BLADE checks if each analytical step—conceptual variable selection, admissible transformations, model family, hyper-parameters—corresponds to one of multiple expert-produced solutions (to account for alternatives), which however still limits the amount by which the agent can transform the task; DiscoveryBench instead scores at the level of the final context, variables and relationship identified, using GPT-4 to judge the semantic match between the agent’s claim and the human-produced reference solutions, without considering how the former was obtained. An overview of BLADE and DiscoveryBench can be found in Figs. 3 and 6. A recent paper (Hsu et al., 2025) has developed a benchmark, WebDS, to tackle the limitations of current benchmarks in evaluating LLM agents on their ability to browse and search the internet over multiple, multimodal, unstructured websites to identify and collect datasets dynamically and perform the full trajectory of web data science activities, where they define a web data science task as one that requires navigation within an environment to acquire new raw information, that is then manipulated. They manually write 870 tasks based on 29 websites covering 10 data-heavy domains; they evaluate both the end results (using automated evaluation metrics where ground-truths are available), as well as the intermediate steps (such as subtask completion, tool usage, data validity, reasoning quality) using an LLM-as-a-judge which gives scores on a scale from 1 to 5, and then validating the protocol with human studies and stability analysis.
@@
From Table 3, we can see that agents have been evaluated on more data science activities than assistants (Table 2), particularly considering goal-oriented activities (with the exception of Business Understanding); however, there is still a lack of evaluations for data management and, to a lesser extent, exploratory activities. Additionally, of all the surveyed works, only BiasBenchmark (Li et al., 2025b), CoTa (Li et al., 2025c), and IDA-Bench (Li et al., 2025a) evaluate agents in a collaborative framework with (simulated) users; whilst simulated users may be an oversimplification of real users interactions, and ignores the employment of domain experts in real-world scenarios, it is a step in the direction of evaluating interactions and their impact on performance.
@@ 5 Challenges and future directions
Our analysis shows that most evaluation works focus either on assistance, looking at isolated tasks that require LLMs to provide an answer on a single-turn basis (without access to tools and under human supervision) or on full automation, where LLMs are wrapped in agents that act autonomously. A few notable exceptions exist (Yu et al., 2019a; Li et al., 2025c; b; a), which primarily rely on other LLMs to simulate human users. While this approach ensures cost-effective evaluation and reproducibility, exploratory tasks may lead the agent to attempt novel solutions, such that the simulated user may be unable to assist it as the most suitable answer may not be within its knowledge base. This can be partly addressed by employing humans to answer such unprecedented queries and progressively enriching the knowledge base. Overall, however, the evaluation ecosystem should be enriched by more fully-fledged “centaur evaluations” (Haupt & Brynjolfsson, 2025) that directly evaluate human-AI collaborations, reorienting AI development towards augmentation rather than substitution and allowing researchers to directly measure human-centred desiderata, such as perceived helpfulness; among the surveyed works, only Li et al. (2025a) goes towards these more complex metrics, by attempting to quantify trade-off between autonomy and reliability/performance.
We found data management and exploratory activities remain mostly uncovered. This is due to 1) the inherent difficulty of scoring exploratory activities, which lack a fixed ground truth, and 2) the complex real-world interactions that certain exploratory activities (such as Business Understanding and Goal Exploration) and data management activities (Acquisition and Simulation) demand. To address this issue, simulated environments where data management and client interaction can occur should be developed, analogous to related developments in scientific research evaluation (Jansen et al., 2024; Cerrato et al., 2025). Such environments would enable the evaluation of agents or assistants that function holistically as data scientists by understanding business requirements, facilitating data collection, and adapting customer requests through data exploration. Simultaneously, realistic evaluation should progress toward end-to-end tasks that do not depend on strict ground truth comparisons or simple activity-specific metrics. Instead, these evaluations should reward insight generation (such as the works in Sec. 4.3) in potentially original ways, thereby properly incentivising systems that fundamentally redefine activities rather than focusing solely on human substitution.
Overall, these gaps in the evaluation landscape make it impossible to obtain a comprehensive characterisation of the performance of LLM-based systems across data science tasks and activities. This obscures potentially strong correlations between specific activities, which may better illuminate the underlying capability space and provide insights on how to best improve these systems going forward, following the traditional path in machine learning where better evaluations in particular domains precede better systems. Thus, we propose the following actions to improve data science evaluation of AI systems:
@@
  • Once gaps in the evaluation landscape are filled, conduct correlation studies on performance across the different data science activities.
@@
Figure 1: Martínez-Plumed et al. (2019): “The DST map, containing the outer circle of exploratory activities, inner circle of CRISP-DM (or goal-directed) activities, and at the core the data management activities.”
@@ Appendix B Appendix: examples of tasks
Figure 5: Li et al. (2025c): The figure shows the “Normal” mode, with the agent being provided all the relevant information and tasked with writing code to address the task, and “Action” mode, where the agent has to take a specific action (in this case, asking for clarification). “Private” refers to tasks requiring the use of bespoke software libraries to which the agent has access to.